Please help me to solve this sumIf tanA=2,tanB=3 then the value of (A+B)

Grade:11

1 Answers

CHAKRAVARTHI
6 years ago
SEE GIVEN,
tan A=2
     A=tan-1(2)..............(1)  [take it as equation no.1]
AND,
2tan B=3,
       B=tan-1(3).............(2)  [take it as equation no.2]
 
hence ,(1)+(2)= A+B=tan-1(2)+tan-1(3)